# Spinal open books and symplectic fillings with exotic fibers

### v1

@@ 6.1. Fillings of planar spinal open books
Thus $(T_{-}(n),\xi_{1})$ is not strongly fillable if $n<-4$ and there is a unique Stein filling for $\xi_{1}$ if $n\geq-4$. ∎
@@
Thus $\phi_{-5}$ is a product of $\phi_{-2}$ and $\phi_{-3}$. According to Theorem 1.8, however, $(P,\phi_{n},\Sigma)$ is Stein fillable for $n=-2,-3$, but not strongly fillable for $n=-5$. ∎
@@
Consider the contact $3$-manifold $(M,\xi)$ supported by the planar spinal open book with two paper components with annulus fibers and monodromy interchanging boundary components, and a single spine component with $\Sigma_{2,2}$ vertebra. The Stein fillings of this manifold are either bordered Lefschetz fibrations with $\Sigma_{1,2}$ base and annulus fibers with no singular fibers, or a PANLF with annulus base, annulus fibers, and eight singular fibers.
@@
In the case $B=\Sigma_{0,2}$, by Riemann–Hurwitz formula, there are four branch points. By Theorem C, the corresponding fillings are PANLFs with annulus fibers and annulus base, and their completions have 4 exotic fibers. The PANLF with annulus fibers and annulus base, whose completion has $4$ exotic fibers, and no singular fibers, induces a spinal open book on its boundary with two page components with annulus fibers, such that monodromy on one of the page components interchanges boundary components, while on the other page component the monodromy is composition of three boundary interchanges, which is equivalent to one boundary exchange composed with $\tau_{\gamma}^{-8}$ by Proposition 6.2. Notice that the monodromy representation must be trivial in this case. To agree with the spinal open book monodromy as in the statement of the theorem, the PANLF filling must also contain 8 singular fibers, each of which has $\gamma$ as its vanishing cycle. This gives the unique minimal filling of $(M,\xi)$ that contains exotic fibers.

### v2

@@ 6.1. Fillings of planar spinal open books
Thus $(T_{-}(n),\xi_{1})$ is not strongly fillable if n<4n<-4 and there is a unique Stein filling for $\xi_{1}$ if $n\geq-4$. ∎
@@
Thus $\phi_{-5}$ is a product of $\phi_{-2}$ and $\phi_{-3}$. According to Theorem 1.8, however, $(P,\phi_{n},\Sigma)$ is Stein fillable for n=2,3n=-2,-3, but not strongly fillable for n=5n=-5. ∎
@@
Consider the contact $3$-manifold $(M,\xi)$ supported by the planar spinal open book with two paper components with annulus fibers and monodromy interchanging boundary components, and a single spine component with $\Sigma_{2,2}$ vertebra. The Stein fillings of this manifold are either bordered Lefschetz fibrations with $\Sigma_{1,2}$ base and annulus fibers with no singular fibers, or a PANLF with annulus base, annulus fibers, and four singular fibers.
@@
In the case $B=\Sigma_{0,2}$, by Riemann–Hurwitz formula, there are four branch points. By Theorem C, the corresponding fillings are PANLFs with annulus fibers and annulus base, and their completions have 4 exotic fibers. The PANLF with annulus fibers and annulus base, whose completion has $4$ exotic fibers, and no singular fibers, induces a spinal open book on its boundary with two page components with annulus fibers, such that monodromy on one of the page components interchanges boundary components, while on the other page component the monodromy is composition of three boundary interchanges, which is equivalent to one boundary exchange composed with $\tau_{\gamma}^{-4}$ by Proposition 6.2. Notice that the monodromy representation must be trivial in this case. To agree with the spinal open book monodromy as in the statement of the theorem, the PANLF filling must also contain 4 singular fibers, each of which has $\gamma$ as its vanishing cycle. This gives the unique minimal filling of $(M,\xi)$ that contains exotic fibers.